OpenStudy (aaronandyson):

Divide 8 into two parts such that the sum of their reciprocals is 8/15 @UnkleRhaukus

OpenStudy (unklerhaukus):

sum of their reciprocals*?

OpenStudy (aaronandyson):

Yes.

OpenStudy (unklerhaukus):

8 = a+b (1/a) + (1/b) = 8/15

OpenStudy (unklerhaukus):

combine these terms: (1/a) + (1/b) like you would any of other fractions

OpenStudy (aaronandyson):

They have given a hint let the two parts be 8,8-x

OpenStudy (aaronandyson):

Like \[\frac{ 1 }{ 8 } + \frac{ 1 }{ 8 - x } = \frac{ 8 }{ 15 }\] But how??

zepdrix (zepdrix):

Hmm that seems a lot more straight forward than the way I did it :D lol cool problem though :)

OpenStudy (unklerhaukus):

i don't think the method i am suggesting makes use of the hint

OpenStudy (aaronandyson):

So,what next @zepdrix @UnkleRhaukus ?

OpenStudy (unklerhaukus):

(1/a) + (1/b) = (b/ab) + (a/ab) = ( . . . ) /ab

OpenStudy (aaronandyson):

Can you use the hint please???

OpenStudy (unklerhaukus):

The hint does not make sense. if 8 = 8 + (8-x), then x = 8, the second part is empty, and you haven't divided the 8 at all

OpenStudy (aaronandyson):

But,They just can't give a USELESS hint in all probability.

zepdrix (zepdrix):

Rhauk, I think this is what they had intended by the hint, (this is actually the path I took). 8= a + b -> a=(8-b) Substitute that for all the a in the other equation.

zepdrix (zepdrix):

not all the a, the a.. blah

zepdrix (zepdrix):

the one single a lol

OpenStudy (aaronandyson):

????? I'm confused now :(

zepdrix (zepdrix):

They were likely calling the who numbers x and y. So 8 = x + y. y = 8-x. 8 = x + (8-x)

zepdrix (zepdrix):

the two* wow my typos t.t

OpenStudy (unklerhaukus):

I found the question much easier without the hint.

OpenStudy (aaronandyson):

???????????????????????????????????????????

OpenStudy (unklerhaukus):

If you want to use the hint, don't follow my method.

OpenStudy (aaronandyson):

I'm just lost now :(

OpenStudy (unklerhaukus):

I don't understand the method with the hint. —soory

OpenStudy (aaronandyson):

I got x^2 - 8x - 15 = 0 @zepdrix

OpenStudy (aaronandyson):

5 & 3?

OpenStudy (unklerhaukus):

i think a sign is wrong in the equation " x^2 - 8x - 15 = 0 " ,

OpenStudy (aaronandyson):

+15* My typos X_X

OpenStudy (unklerhaukus):

now check you answer, does 8 = 5+3 ? does 1/5+1/3 = 3/(3*5) + 5/(3*5)= 8/15 ?

OpenStudy (unklerhaukus):

yes/no, yes/no

OpenStudy (aaronandyson):

Idk.

OpenStudy (unklerhaukus):

DOES 8 = 5 + 3 ? count on your finders if you have to

OpenStudy (unklerhaukus):